4. A block of mass 3.0 kg slides with uniform velocity down a plane inclined 25o with the horizontal. If the angle of inclination is increased to 40o , what will be the acceleration of the block? 


1
Expert's answer
2021-09-21T06:34:18-0400

Given:

"\theta_1=25^{\circ},\quad \theta_2=40^{\circ}"


The Newton's second law gives

"a=g(\sin\theta-\mu\cos\theta)"

The coefficient of kinetic friction is given by

"\mu=\tan\theta_1"

Hence, the acceleration of the body

"a=g(\sin\theta_2-\tan\theta_1\cos\theta_2)"

"a=9.8\:{\rm m/s^2}(\sin40^{\circ}-\tan25^{\circ}\cos40^{\circ})=2.8\:\rm m/s^2"